# Conceptualizing Uncertainty

### v1

@@ Abstract
Uncertainty in machine learning refers to the degree of confidence or lack thereof in a model’s predictions. While uncertainty quantification methods exist, explanations of uncertainty, especially in high-dimensional settings, remain an open challenge. Existing work focuses on feature attribution approaches which are restricted to local explanations. Understanding uncertainty, its origins, and characteristics on a global scale is crucial for enhancing interpretability and trust in a model’s predictions. In this work, we propose to explain the uncertainty in high-dimensional data classification settings by means of concept activation vectors which give rise to local and global explanations of uncertainty. We demonstrate the utility of the generated explanations by leveraging them to refine and improve our model.  [^1] Code will be freely available here (upon acceptance): https://github.com/robertsi20/Conceptualizing-Uncertainty.
@@ 1 Introduction
While advances in deep learning in the last years have led to impressive performance in many domains, such models are not always reliable, particularly when it comes to generalizing to new environments or adversarial attacks. To improve on that, numerous methods have been developed in the field of explainable artificial intelligence (xAI) [5] to provide insights into model behavior and facilitate actionable modifications. However, the majority of methods focus on explaining model predictions, which can help understand misclassifications but do not explicitly address predictive uncertainty(See Figure 1). Understanding uncertainty is crucial for detecting potential model weaknesses, particularly in dynamic environments.
@@
Since uncertainty quantification is useful in various applications, including active learning [20], classification with rejects [17], adversarial example detection [26], and reinforcement learning [24], a significant body of work aims to improve the quantification of predictive uncertainty using Bayesian deep learning (BDL) and approximations thereof [15, 9, 14]. In contrast, the literature on understanding the sources of uncertainty for a given model via explanations is limited, focusing on methods for feature attribution [28, 27] (see section 2.4 for more related work). While these approaches do provide valuable insights for individual input data points, they are restricted to local explanations and do not provide high-level insights on a global dataset scale, e.g., for image datasets. Also, another limitation of feature attribution methods, known from the prediction attribution setting and prevalent in the image domain, is that these point to which part of the input the model considers as important, but not why [13]. Here, methods aiming to automatically extract concept activation vectors (CAVs) [11, 1] and attributing them back to the input do provide both, information on what the model seems to perceive in a part of the input, and global dataset information, both by inspecting the learned concepts.
@@
In this work, (i) we propose a pipeline to enable concept-based explanations for predictive uncertainty, providing both local and global explanations for sources of uncertainty, and (ii) we demonstrate the correctness and usefulness of the learned concepts as well as the potential to perform actionable interventions based on these in a series of downstream tasks, including the automatic detection of different types of novelty in new environments, improving uncertainty-based rejections and detecting gender bias in language models.
We proceed by describing foundations for our contributions, including related work on uncertainty explanations. This is followed by our proposed pipeline in section 3, experiments in 4, a discussion on limitations in 5 and a conclusion.
Figure 1: A White Wolf classified as Samoyed by an ImageWoof-adapted ResNet50 classifier. We generated the concept attribution map explaining the prediction (left) versus explaining the uncertainty (right) with our proposed pipeline. Correspondingly, we display three important concepts for this image using each explanation approach. The prediction explanation suggests the neural net arrived at its decision due to the detected concepts of a Samoyed’s ears, face, and legs, while the uncertainty explanation offers that the model is uncertain about its prediction because of the detected concept of a wolf’s head, fur, and legs.
@@ 2.1 Setup and Problem Formulation
Given a trained classification model $\mathbf{M}$, e.g., a deep convolutional network, and a dataset of $n$ data points $\mathbf{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{n}\}$, usually not seen during training. In contrast to a vast amount of literature focusing on explaining predictions of such models [5], we aim for explaining their predictive uncertainty, thereby aiming to understand its sources. Our work differs from existing recent research on that topic, e.g., [5], in that we aim for global explanations.
For the following Bayesian formalization, we also require the training data set $\mathcal{D}$ of $\mathbf{M}$, which, however, is not used by our method.
@@ 2.2 Quantifying Uncertainty
Predictive uncertainty is typically defined over the predictive distribution $p(y|\mathbf{x},\mathcal{D})$ [15, 9], which in our case is over possible labels $y$. In a Bayesian setting, where the parameters $\boldsymbol{\theta}$ of our classification model $\mathbf{M}$ are random variables, $p(y|\mathbf{x},\mathcal{D})=\mathbb{E}_{p(\boldsymbol{\theta}|\mathcal{D})}[p(y|\mathbf{x},\boldsymbol{\theta})]$ requires computing the expectation $\mathbb{E}$ over the posterior $p(\boldsymbol{\theta}|\mathcal{D})$, which in practice is usually intractable. Accordingly, various methods for posterior approximation have been introduced such as ensemble methods [21], Monte Carlo (MC) dropout [14] and Laplace Approximations [6].
Different measures for predictive uncertainty are defined in literature [15, 9]. We summarize the ones we are using in our experiments in the following.
@@
In order to approximate the above measures, we utilize MC dropout to collect a set predictions $\{p(y|\mathbf{x},\boldsymbol{\theta}_{i})\}_{i=1}^{N}$ and approximate the posterior predictive $p(y|\mathbf{x},\mathcal{D})=\mathbb{E}_{p(\boldsymbol{\theta}|\mathcal{D})}[p(y|\mathbf{x},\boldsymbol{\theta})]\approx\frac{1}{N}\sum_{i}^{N}p(y|\mathbf{x},\boldsymbol{\theta}_{i}).$ We refer to $\hat{u}_{t},\hat{u}_{a},\hat{u}_{e}$ when utilizing this approximation in $u_{t},u_{a},u_{e}$, respectively.
@@ 2.3 Concept Activation Vectors
Concept Activation Vectors (CAVs) aim for human interpretability with respect to understanding black-box model predictions [19, 11, 16]. Recent approaches based on Nonnegative Matrix Factorization (NMF) [13, 18] have been shown to demonstrate superior qualitative and quantitative properties of the resulting concepts[11]. They work by embedding patches of inputs into a non-negative activation space of a pre-trained model and applying NMF to decompose the embedded data matrix $\mathbf{A}$ into a product of non-negative matrices $\mathbf{W}$ and $\mathbf{V}$, solved by reconstructing $\mathbf{A}$, i.e.,  $(\mathbf{W},\mathbf{V})=\arg\min_{\mathbf{W}\geq 0,\mathbf{V}\geq 0}\|\mathbf{A}-\mathbf{W}\mathbf{V}^{\top}\|^{2}_{F}$. The decomposition yields: $\mathbf{V}$ the dictionary of concepts (or concept bank) and $\mathbf{W}$ a reduced representation of $\mathbf{A}$ according to the basis $\mathbf{V}$. To attribute importance, they make use of a sensitivity analysis technique known as total Sobol Indices, which captures the effects of a concept along with its interactions on the model’s output by considering the variance fluctuations by perturbing $\mathbf{W}$.
@@ 2.4 Related Work on Explaining Uncertainty
While there does exist a plethora of xAI methods for explaining the prediction of classification models including several local and global approaches [5], methods for explaining the source of uncertainty have only developed in the last years. So far, these have focused on local feature attribution explanations, including explaining uncertainty with shapley values [28], with gradient-based methods [27], with counterfactuals [2] and by taking second order effects into account [4]. In contrast, we aim for explanations beyond feature attribution, that provide also global explanations of uncertainty, which enable an overview of characteristics of uncertain predictions. Also, concerning image datasets, rather small ones consisting of MNIST and CIFAR are used in [28, 27, 2]. Only [4] is applied to CelebA, containing larger images. We apply our approach to images of ImageNet.
Another, less related line of work, aims for uncertainty sets, dealing with the problem of uncertainty of explanations [22].
@@ 3 Proposed Pipeline
We propose to explain predictive uncertainty by means of CAVs computed on a local level – for each data point – or aggregated to obtain a global explanation. In Figure 2, we illustrate our proposed pipeline which aims to characterize and explain uncertainty using extracted concepts using high-dimensional data $\mathcal{D}$. Given also a classification model $\mathbf{M}$, and an uncertainty measure $u$, e.g., $u\coloneq\hat{u}_{t}$, we arrange our pipeline into 4 steps:
Step 1 - We use the model $\mathbf{M}$ to classify the inputs and compute $u(\mathbf{x})$ for each data point, by leveraging an approximation technique such as MC Dropout.
Step 2 - To apply concept activations, we need to specify a probabilistic classification task. In our setup, a canonical choice is the grouping induced by uncertain (UNC) and certain (CER) samples. To obtain a reasonable decision boundary, we observe that $u$ will map CER to small numbers and UNC to large numbers. We therefore expect that if applied to all data points, $\{u(\mathbf{x}_{i})\}_{i=1}^{n}$ can be described by a mixture model with two components. For simplicity, we assume a Gaussian Mixture Model (GMM) with two components which we train on $\{u(\mathbf{x}_{i})\}_{i=1}^{n}$. By our assumption we expect the component with the larger mean to correspond to the UNC samples. We thus obtain the classification model $f$ by considering the conditional probability which takes on a sigmoid shape. To counteract potential class imbalance, i.e., an imbalance between CER and UNC in the training data, we set the weights of the mixture component to equal size after training.
@@ 4 Experiments
Since uncertainty arises from multiple sources, establishing a ground truth for experimentation can be challenging. Therefore, we aim to demonstrate the validity and usefulness of our uncertainty explanations by integrating them into a classification with reject options setting. Additionally, we illustrate how our concepts capture different sources of uncertainty, aiding human decision-makers in constructing effective re-training sets. Finally, we show that our concepts can reveal potential biases with respect to sensitive groups in a downstream task.

### v2

@@ Abstract
Uncertainty in machine learning refers to the degree of confidence or lack thereof in a model’s predictions. While uncertainty quantification methods exist, explanations of uncertainty, especially in high-dimensional settings, remain an open challenge. Existing work focuses on feature attribution approaches which are restricted to local explanations. Understanding uncertainty, its origins, and characteristics on a global scale is crucial for enhancing interpretability and trust in a model’s predictions. In this work, we propose to explain the uncertainty in high-dimensional data classification settings by means of concept activation vectors, which give rise to local and global explanations of uncertainty. We demonstrate the utility of the generated explanations by leveraging them to refine and improve our model.  [^1] Code is freely available here: https://github.com/robertsi20/Conceptualizing-Uncertainty.
@@ 1 Introduction
While advances in deep learning in recent years have led to impressive performance in many domains, such models are not always reliable and pose risks in real-world applications, especially when exposed to dynamic environments. As such, numerous methods have been developed in the field of explainable artificial intelligence (xAI) [5] to provide insights into model behavior and facilitate actionable modifications. However, most methods focus on explaining model predictions, which does not explicitly address predictive uncertainty (see Figure 1). Understanding sources of uncertainty is crucial for detecting potential model weaknesses and data flaws and, additionally, provides means of meaningful downstream actions [19], aimed at increasing trust and reliability.
Understanding uncertainty and its sources requires 3 main steps: 1. localizing it, 2. assigning a degree of uncertainty, and 3. finding its origin. As such, Uncertainty Quantification (UQ) methods emerged as a tool and have proven useful in various applications, including active learning [22], classification with rejects [17], adversarial example detection [34], reinforcement learning [28], and separating sources of uncertainty [19]. Therefore, a significant body of work aims to improve the quantification of predictive uncertainty using techniques such as Bayesian deep learning (BDL) and approximations thereof [15, 9, 14]. However, existing methods of UQ assign values of certainty without providing human-interpretable insights about what may be causing it [15].
@@
For this reason, some attempts have been made to explain uncertainty with state-of-the-art xAI techniques, including feature attributions [36, 35] and counterfactuals [2]. Although these approaches provide valuable information for individual data points, they are restricted to local explanations and do not provide high-level information on a global dataset scale, such as in image datasets. Additionally, feature attribution methods are limited, because they point only to which part of the input the model considers as important, but do not explain why [13]. In this regard, methods aiming to automatically extract concept activation vectors (CAVs) [11, 1] and attribute them back to the input provide both information on what the model seems to perceive in a part of the input and global dataset information, by inspecting the learned concepts.
Figure 1: We display concept-based explanations of a model prediction from literature [13] (left) and our proposed concept-based explanations of predictive uncertainty (right) for the same input. Each explanation contains three most important concepts, each visualized by the 6 most activating patches from the training set. The prediction explanation suggests the neural net arrived at its decision primarily due to the detected concept 1 of the background, while the uncertainty explanation offers that the model is uncertain about its prediction largely because of the detected concept i of the noise.
@@
In this work, (i) we propose a novel pipeline to enable concept-based explanations for predictive uncertainty, providing both local and global explanations for sources of uncertainty in a human-interpretable way, and (ii) we demonstrate the potential to perform actionable interventions based on the learned concepts in a series of downstream tasks, including the automatic detection of different types of uncertainty in new environments, interpretable uncertainty-based rejections and detecting gender bias in language models, thereby showing their usefulness.
This paper is organized as follows: Section 2 provides background information needed to understand our proposed pipeline as well as related work. In Section 3, we provide the details of our pipeline. Section 4 presents the experiments where we evaluate the pipeline’s effectiveness. We finish our work with a discussion on limitations in Section 5 and a conclusion.
@@ 2.1 Setup and Problem Formulation
Given a trained classification model $\mathbf{M}$, e.g., a deep convolutional network, and a dataset of $n$ data points $\mathbf{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{n}\}$, usually not seen during training. In contrast to a vast amount of literature focusing on explaining predictions of such models [5], we aim to explain their predictive uncertainty, thereby aiming to understand its sources. Our work differs from recent research on that topic, e.g., [5], in that we aim for human-interpretable concept-based explanations.
For the following UQ, we proceed with a Bayesian formalization due to its precedence [15, 19] in the literature. As such, we also require the training data set $\mathcal{D}$ of $\mathbf{M}$, which, however, our method does not use.  [^2] Other notable frameworks include Conformal Prediction [32] and Frequentist approaches [15] and could also be used with our pipeline.
@@ 2.2 Quantifying Uncertainty
A popular way to define predictive uncertainty is over the predictive distribution $p(y|\mathbf{x},\mathcal{D})$ [15, 9], which in our case is over possible labels $y$.
In our Bayesian setting, where the parameters $\boldsymbol{\theta}$ of our classification model $\mathbf{M}$ are random variables, $p(y|\mathbf{x},\mathcal{D})=\mathbb{E}_{p(\boldsymbol{\theta}|\mathcal{D})}[p(y|\mathbf{x},\boldsymbol{\theta})]$ requires computing the expectation $\mathbb{E}$ over the posterior $p(\boldsymbol{\theta}|\mathcal{D})$, which is usually intractable. Accordingly, various methods for its approximation have been introduced [15], such as Variational Inference based approaches like Monte Carlo (MC) dropout [14], sampling based methods [37], Laplace Approximations [6] and ensembles [23].
Different measures for predictive uncertainty are defined in the literature [15, 9]. We summarize and use the measures in the following due to their prominence in the literature [36, 35], but our explanation approach can be applied with any measure. We can also compute metrics that quantify uncertainty sources into their aleatoric and epistemic components. Since this is a classical way of explaining sources of uncertainty, we include them for comparison to our method.
@@
To approximate the above measures, we utilize MC dropout to collect a set of predictions $\{p(y|\mathbf{x},\boldsymbol{\theta}_{i})\}_{i=1}^{N}$ and approximate the posterior predictive $p(y|\mathbf{x},\mathcal{D})=\mathbb{E}_{p(\boldsymbol{\theta}|\mathcal{D})}[p(y|\mathbf{x},\boldsymbol{\theta})]\approx\frac{1}{N}\sum_{i}^{N}p(y|\mathbf{x},\boldsymbol{\theta}_{i}).$ We refer to $\hat{u}_{t},\hat{u}_{a},\hat{u}_{e}$ when utilizing this approximation in $u_{t},u_{a},u_{e}$, respectively.
@@ 2.3 Concept Activation Vectors
Concept Activation Vectors (CAVs) aim for human interpretability with respect to understanding black-box model predictions [21, 11, 16]. These can be categorized into two classes, concept bottleneck models which enforce the use of concepts during training and post-hoc methods that are applied after training and provide additional information as compared to saliency maps [13]. In the present work, we focus on such post-hoc approaches, since concept bottleneck models usually require concept labels (apart of some notable exceptions [27]) and we are interested in concepts that explain uncertainty. Recent approaches based on Nonnegative Matrix Factorization (NMF) [13, 29, 20] have been shown to demonstrate superior qualitative and quantitative properties of the resulting concepts [11]. Here, (parts of) the input are typically embedded into a non-negative activation space of a pre-trained model and NMF decomposes the embedded data matrix $\mathbf{A}$ into a product of non-negative matrices $\mathbf{W}$ and $\mathbf{V}$, solved by reconstructing $\mathbf{A}$, i.e., $(\mathbf{W},\mathbf{V})=\arg\min_{\mathbf{W}\geq 0,\mathbf{V}\geq 0}\|\mathbf{A}-\mathbf{W}\mathbf{V}^{\top}\|^{2}_{F}$. The decomposition yields: $\mathbf{V}$ the dictionary of concepts (or concept bank) and $\mathbf{W}$ a reduced representation of $\mathbf{A}$ according to the basis $\mathbf{V}$. To attribute importance, the authors of [13] make use of a sensitivity analysis technique known as total Sobol Indices, which captures the effects of a concept along with its interactions on the model’s output by considering the variance fluctuations by perturbing $\mathbf{W}$. The contribution of concept $i$ is then defined by: $S_{i}^{T}=\mathbb{E}_{\mathbf{M}_{\sim i}}\left(\mathbb{V}_{\mathbf{M}_{i}}(h((\mathbf{W}\circ\mathbf{M})\mathbf{V}^{\top})|\mathbf{M}_{\sim i})\right)\backslash\mathbb{V}\left(h((\mathbf{W}\circ\mathbf{M})\mathbf{V}^{\top})\right)$, where $h$ is the model function mapping embeddings $\mathbf{A}$ to the model’s output, $\mathbf{M}$ are uniformly and i.i.d. stochastic masks in $[0,1]^{r}$ with $r$ concepts, $\circ$ the Hadamard product, $\sim$ the complementary function [13].
@@ 2.4 Related Work on Explaining Uncertainty
While a plethora of xAI methods exists for explaining the prediction of classification models, including several local and global approaches [5], methods for explaining the source of uncertainty have only developed recently. Mostly, these have focused on local feature attribution explanations, including explaining uncertainty with shapley values [36], with gradient-based methods [35], with counterfactuals [2] and by taking second-order effects into account [4]. In contrast, we aim for explanations beyond feature attribution that also provide global explanations of uncertainty, which enable an overview of uncertain prediction characteristics. Also, concerning image datasets, rather small ones consisting of MNIST and CIFAR are used in [36, 35, 2]. Only [4] is applied to CelebA, containing larger images. We apply our approach to images of ImageNet.
A few methods have been proposed for explaining uncertainty on a global scale. These include [26], which utilizes dimensionality reduction and an adaptation of [31] to visualize uncertainty patterns and [38], using conformal prediction.
Another, less related line of work, aims to provide uncertainty estimates of explanations [30], e.g., by utilizing uncertainty sets [24]. We, in contrast, aim to explain the uncertainty of the classifier, not the uncertainty of an explanation.
@@ 3 Proposed Pipeline
We propose to explain predictive uncertainty by means of CAVs computed on a local level – for each data point – or aggregated to obtain a global explanation. In Figure 2, we illustrate our proposed pipeline, which aims to characterize and explain uncertainty using extracted concepts from high-dimensional data $\mathbf{X}$, by grouping the predictions into certain and uncertain ones and extracting concepts from these. More specifically, given such data $\mathbf{X}$, a classification model $\mathbf{M}$, and an uncertainty measure $u$, e.g., $u\coloneq\hat{u}_{t}$, we arrange our pipeline into 4 steps:
Step 1 - We use the model $\mathbf{M}$ to classify the inputs and compute $u(\mathbf{x}_{i})$ for each data point, by leveraging an approximation technique such as MC Dropout.
Step 2 - In order to group $u(\mathbf{x}_{i})$ into uncertain (UNC) and certain (CER) samples, we specify a probabilistic classification task $f:u(\mathbf{x})\mapsto[0,1]$, such that $f(u(\mathbf{x}))<0.5$ corresponds to CER samples. We therefore expect that if applied to all data points, $\{u(\mathbf{x}_{i})\}_{i=1}^{n}$ can be described by a mixture model with two components. For simplicity, we assume a Gaussian Mixture Model (GMM) with two components, which we train on $\{u(\mathbf{x}_{i})\}_{i=1}^{n}$. By our assumption, we expect the component with the larger mean to correspond to the UNC samples. We thus obtain the classification model $f$ by considering the conditional probability, which takes on a sigmoid shape.
@@ 4 Experiments
Since uncertainty arises from multiple sources, establishing a ground truth for experimentation can be challenging. Therefore, we aim to demonstrate the validity and usefulness of our uncertainty explanations by illustrating how our concepts capture different sources of uncertainty, aiding human decision-makers in constructing effective re-training sets. Additionally, we integrate them into a classification with reject options setting. Finally, we show that our concepts can reveal potential biases concerning sensitive groups in a downstream task.
Figure 3: Setup of Experiment 4.1: Points with high uncertainty (Phase 1) are used for concept extraction (Phase2). Concepts are inspected manually for unstructured noise (concept 2 is identified here). Uncertain points are grouped according to the activation of the noise concept automatically (Phase 3).
Hyperparameters For our proposed approach, we utilize the following choice of hyperparemeters: patch size: same as in the craft paper for vision? for language?; number of concepts - is there a reason there is 10 per cert and uncert in experiment 1, and 55 for cert and 35 for uncert in experiment 2? Both have 10 classes? layer - penultimate, following [11]; backbone network - for vision tasks we use ResNet-50 and for language a BERT model
4.1 Distinguishing Sources for Uncertainty
Figure 4: Experiment 4.1 results for Motion Blur ($\uparrow$): When grouping uncertain data for re-training, we list the percentage of useful data among the selected ones (y-axis) for the different percentages of selected data (x-axis).
Specifically, the goal is to separate sources of uncertainty, and we show its benefit by building a representative re-training set in a setting where a deployed model $\mathbf{M}$ is exposed to two noise types: 1. Structured noise, such as the introduction of novel classes not seen during training, and 2. Unstructured noise, such as random distortions like blurring. We assume that the unstructured noise has rendered the image unusable, such that, in the context of re-training, no meaningful signal can be derived from the affected inputs.
For more clarity, we include Figure 3. We begin by applying Steps 1 & 2 from our pipeline to $\mathbf{M}$ to obtain the uncertain inputs. Then, with Step 3, we extract their concepts. Since concepts provide a human understanding of the source of uncertainty, a practitioner visually inspects the concepts by examining the patches most activated by each concept in order to identify those associated with the unstructured noise. In Phase 2 of Figure 3, the middle concept describes a blur. Now, using our explanations, we can automatically locate the inputs that most heavily activate this concept and exclude them from our re-training set.
More precisely, we consider the setting where new data samples $\mathbf{X}$ are available and our explanation pipeline is applied to provide concept banks $\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}$ of 10 concepts each, according local importances $e_{l}(\mathbf{x}_{i}),\forall\mathbf{x}_{i}\in\mathbf{X}$ and global ones $e_{\text{CER}},e_{\text{UNC}}$. We aim to evaluate how well $\mathbf{V}_{\text{UNC}}$ captures different noise types, and how well unstructured noise samples $\mathbf{x}_{i}$ can be filtered out using $e_{l}(\mathbf{x}_{i})$ and $\mathbf{W}_{i}$ of $\mathbf{V}_{\text{UNC}}$ concepts. For this purpose, we assume that we can determine the set $nc$ that corresponds to unstructured noise concepts in $\mathbf{V}_{\text{UNC}}$ (through visual inspection). Then, for each data point $\mathbf{x}_{i}$, we sum the local importance $e_{l,nc}(\mathbf{x}_{i})$ or NMF activations $\mathbf{W}_{i,nc}$ and filter out according to the highest values, corresponding to the amount of presence of these concepts. We refer to these strategies as Ours (Importance) when using $e_{l,nc}(\mathbf{x}_{i})$, and to and Ours (NMF) for $\mathbf{W}_{i,nc}$.
[…]
Table 1: We report the average AUC score ()($\uparrow$)( ↑ ) over 20 runs for various types of noise patterns. Our proposed methods perform better than the baselines. We test our methods against Gaussian Blurring (G Blur), Salt and Pepper Noise (S and P), Gaussian Noise (G Noise), Motion Blurring (M Blur), Radial Blurring (R Blur), and Wave Noise(Wave).
Method Total Aleatoric Epistemic Ours (Imp) Ours (NMF)
[…]
The results in Table 1 indicate that our method outperforms uncertainty-based measures for this task. Thereby, using $e_{l}(g(\mathbf{x}_{i}))$ achieves the second-best performance (underlined), while leveraging $\mathbf{W}_{i}$ yields the best results (bolded), with the highest average AUC score. Inspecting Figure 4, we can see that our method using the NMF coefficients (purple) consistently recommends images that would be beneficial to the re-training set while abstaining from recommending the unusable images until they must be chosen. We indicate the true unusable image percentage by the vertical line on the figure. In the ideal case, a method would not recommend any unusable images until it reaches the vertical line. Our method provides an explainable way to automatically select images that contain meaningful structure for domain adaptation.